OpenStudy (steve816):

Last question of my math homework! Finally it's been a long time! Help me and I'm done with math!

OpenStudy (steve816):

Solve please \[\sin(2 \theta) - \cos \theta - 2 \sin \theta + 1 = 0\]

OpenStudy (mathmale):

Steve: Didn't hear back from you regarding our discussion of your previously posted question. How did that turn out?

OpenStudy (steve816):

I figured it out and was right! I substituted the x back with 2 theta and solved.

OpenStudy (steve816):

Thanks for the help mathmale :)

OpenStudy (mathmale):

All right. What are your thoughts regarding this new post? (You're welcome.)

OpenStudy (steve816):

This time, I think I can use the double angle formula since there is both sin and cos in the equation?

OpenStudy (mathmale):

Chances are that this will work. What have you done so far?

OpenStudy (steve816):

\[2\sin \theta \cos \theta - \cos \theta - 2\sin \theta + 1 = 0\]

OpenStudy (mathmale):

I'm trying to factor this expression. Would you try that on your own, please? Is there a common factor in your equation?

OpenStudy (steve816):

I think you can distribute the 2sin theta, so now, I have\[2\sin \theta (\cos \theta - 1) - \cos \theta + 1 = 1\]

OpenStudy (steve816):

Oops, I typed that wrong, I meant = 0

OpenStudy (mathmale):

What do you mean by "distribute the 2 sin theta"?

OpenStudy (steve816):

Sorry, factor out the 2 sin theta

OpenStudy (mathmale):

Try it. I've been looking to factor out cos theta, by the way.

OpenStudy (steve816):

2sinθ(cosθ−1)−cosθ+1=0

OpenStudy (mathmale):

Your format looks very similar to mine. Any idea of what to do next? Again I ask you whether your expression has a common factor.

OpenStudy (steve816):

Hmm, I realize I can cancel out the cosθ−1

OpenStudy (mathmale):

with some manupulation you can make cos theta - 1 into 1 - cos theta.

OpenStudy (mathmale):

Not "cancel out," Steve, but "factor out." If you drop the factor you've cancelled out, you 'll lose a root or two.

OpenStudy (steve816):

\[2\sin \theta (\cos \theta - 1) = \cos \theta - 1\] \[2\sin \theta = \frac{ \cos \theta - 1 }{ \cos \theta - 1 }\] \[2\sin \theta = 1\] \[\sin \theta = \frac{ 1 }{ 2 }\]

OpenStudy (steve816):

I worked out the problem and that is what I got...

OpenStudy (steve816):

Not sure if it's right though

OpenStudy (mathmale):

I've worked thru the same problem in a slightly different way and have come up with 3 solutions on the interval [0,pi]. How would you go about checking the validity of your one solution?

OpenStudy (mathmale):

When solving an equation and coming up with a solution, never assume the solution is correct until after you've substituted it back into the original equation and find the equation to be true.

OpenStudy (steve816):

Oh okay.

OpenStudy (steve816):

So the two solutions I got are pi/6 and 5pi/6 How did you get 3 solutions?

OpenStudy (mathmale):

2sinθ(cosθ−1)−cosθ+1=0 can be re-written as \[2\sin \theta (\cos \theta-1)-( ?? ) = 0\]

OpenStudy (mathmale):

Steve, I asked whether you could factor the equation above. Have you been able to do that? If so, you'll end up with 2 factors, each of which you need to set = to zero separately and solve for theta. This is the procedure that got me 3 solutions on [0,pi].

OpenStudy (steve816):

ohhhh I see what you mean finally!

OpenStudy (mathmale):

;)

OpenStudy (steve816):

Sorry to fail you, but I can't figure out the factors :(

OpenStudy (mathmale):

We have \[2\sin \theta(\cos \theta - 1)-\cos \theta +1 = 0\]

OpenStudy (mathmale):

and the 2nd half of this equation can be re-written to match that \[\cos \theta -1\]

OpenStudy (mathmale):

factor on the left::

OpenStudy (steve816):

ohhhh, i am so stupid I'm sorry, I got this!

OpenStudy (mathmale):

\[2\sin \theta(\cos \theta - 1)-(\cos \theta -1)=0\]

OpenStudy (mathmale):

OK with this?

OpenStudy (steve816):

\[\cos \theta - 1(2\sin \theta - 1) = 0\]

OpenStudy (mathmale):

Can you now factor this latest equation?

OpenStudy (mathmale):

You can try factoring your own equation also. That's the best way, probably, to determine whether your equation will "work" or not. Mind factoring my last equation now?

OpenStudy (mathmale):

My last equation definitely has a common factor. What is that factor?

OpenStudy (steve816):

I finally got it thanks so much for the 30 minutes of your life :)

OpenStudy (mathmale):

It was well spent, Steve. Best of luck to you. Bye for now.